OpenStudy (anonymous):

find a solution to r'(t)=2r(t) where r(t) is a vector valued function in three space @ganeshie8 wheeeeeetttt?

ganeshie8 (ganeshie8):

try \[r(t) = e^{2t}i + e^{2t}j+ e^{2t}k \]

OpenStudy (perl):

interesting, you can get that by integrating componentwise we have the vector valued function r(t) = 2* r ' (t) Let r(t) = < f(t) , g(t) , h(t) > So r ' (t) = < f ' (t) , g ' (t) , h ' (t) > So we have < f ' (t) , g ' (t) , h ' (t) > = 2*< f(t) , g(t) , h(t) > < f ' (t) , g ' (t) , h ' (t) > = < 2*f(t) , 2*g(t) , 2*h(t) > Equating the first components we have f ' (t) = 2*f(t) f ' (t) / f(t) = 2 ln (f(t)) = 2t f(t) =Ce^(2t) similarly g(t) = C*e^(2t) h(t) = C*e^(2t) then it asks for *a* solution, so set C = 1 . but the general solution is r (t) = < C*e^(2t) , C*e^(2t) , C*e^(2t) > you can write this as a linear combination of vectors i,j,k r(t) = C*e^(2t)*i + C*e^(2t)*j + C*e^(2t)* k

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

ok... wait..so am i just allowed to throw in the t next to the 2 and the C as well?

OpenStudy (perl):

you have to plug in a number for C , a nonzero number

OpenStudy (perl):

C = 1 is easiest

OpenStudy (perl):

because it says find 'a' solution, singular

OpenStudy (perl):

i was just showing you how to derive the general solution, fyi

OpenStudy (perl):

oh i left out a step

OpenStudy (perl):

we had f ' (t) = 2*f(t) f ' (t) / f(t) = 2 ln (f(t)) = 2t + c raise both sides to e f(t) =e^(2t + c) by law of exponents f(t) = e^(2t) * e^c now e^constant is still a constant, so let e^c = C f(t) = e^(2t) * C f(t) = C*e^(2t)

OpenStudy (perl):

you always have to justify wherever you place a constant (so i didn't just append a C from out of nowhere).